OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation 2x + 4y = 12. How would I do this? slope = 2; y intercept at (0, 12) slope = 1/2; y intercept at (0, 3) slope = 1/2; y intercept at (0, 4) slope = -1/2; y intercept at (0, 3)

OpenStudy (anonymous):

slope = -1/2; y intercept at (0, 3)

OpenStudy (anonymous):

Start by getting y by itself, what slope do you get?

jimthompson5910 (jim_thompson5910):

2x+4y=12 4y=12-2x 4y=-2x+12 y=(-2x+12)/(4) y=((-2)/(4))x+(12)/(4) y=-(1/2)x+3 So the slope is -1/2 and the y-intercept is (0, 3)